OpenStudy (anonymous):

Consider a game in which players draw one card at a time from a standard 52-card deck. If a player draws a face card (jack, queen, or king) the player is awarded 16 points. Any other card drawn earns the player 3 points. What is the expected value of drawing a car in the game on the first turn? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what's the probability of drawing a face card?

OpenStudy (anonymous):

16/52

jimthompson5910 (jim_thompson5910):

there are 3 face cards (according to their definition and they are: jack, queen, or king)

jimthompson5910 (jim_thompson5910):

so it's NOT 4*4 = 16 out of 52 it's actually 4*3 = 12 out of 52 ----> P(face card) = 12/52

jimthompson5910 (jim_thompson5910):

the probability of drawing anything else is 40/52 I got 40 because 52 - 12 = 40 (ie there are 40 non-face cards)

OpenStudy (anonymous):

Oh okay

jimthompson5910 (jim_thompson5910):

so to find the expected value, you do... Expected Value = E[X] E[X] = P(Face Card)*V(Face Card) + P(Non Face Card)*V(Non Face Card) E[X] = (12/52)*(16) + (40/52)*(3) E[X] = ??? -------------------------------------------- Notes: P(Face Card) = probability of drawing a face card P(Non Face Card) = probability of drawing a non-face card V(Face Card) = the value (prize or penalty) of getting a nonface card V(Non Face Card) = the value (prize or penalty) of getting a nonface card

OpenStudy (anonymous):

I got 6

OpenStudy (anonymous):

Is that right?

jimthompson5910 (jim_thompson5910):

sry just noticed these messages

jimthompson5910 (jim_thompson5910):

yes it's 6

jimthompson5910 (jim_thompson5910):

so on any draw, you expect on average to get roughly 6 points

OpenStudy (anonymous):

Can you help me with another one? I tried to solve it but I didnt get the right answer.

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

Consider a game in which players roll a number cube to determine the number of points earned. If a player rolls a prime number, that many points will be added to the players total. Any other roll will deducted from the players total. What is the expected value of the points earned on a single roll in this game?

jimthompson5910 (jim_thompson5910):

it's a cube, so it has 6 sides

jimthompson5910 (jim_thompson5910):

I'm assuming the 6 faces are: 1, 2, 3, 4, 5, 6

jimthompson5910 (jim_thompson5910):

Of these 6 numbers, the following are prime: 2, 3, 5 Note: 1 is NOT a prime number

jimthompson5910 (jim_thompson5910):

Let's go through each roll 1 by 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Okay

jimthompson5910 (jim_thompson5910):

sry let me redo, one sec

OpenStudy (anonymous):

Sure thing

jimthompson5910 (jim_thompson5910):

Rolling a 1... P(rolling 1) = 1/6 V(rolling 1) = -1 (you lose 1 point if you roll this) multiply this out to get (1/6)*(-1) = -1/6 Note: you lose points because 1 is NOT prime -------------------------------------------------------------- Rolling a 2... P(rolling 2) = 1/6 V(rolling 2) = 2 (you earn 2 points if you roll this) multiply this out to get (1/6)*(2) = 2/6 Note: you earn points because 2 is prime -------------------------------------------------------------- Rolling a 3... P(rolling 3) = 1/6 V(rolling 3) = 3 (you earn 3 points if you roll this) multiply this out to get (1/6)*(3) = 3/6 Note: you earn points because 3 is prime -------------------------------------------------------------- Rolling a 4... P(rolling 4) = 1/6 V(rolling 4) = -4 (you lose 4 points if you roll this) multiply this out to get (1/6)*(-4) = -4/6 Note: you lose points because 4 is NOT prime -------------------------------------------------------------- Rolling a 5... P(rolling 5) = 1/6 V(rolling 5) = 5 (you earn 5 points if you roll this) multiply this out to get (1/6)*(5) = 5/6 Note: you earn points because 5 is prime -------------------------------------------------------------- Rolling a 6... P(rolling 6) = 1/6 V(rolling 6) = -6 (you lose 6 points if you roll this) multiply this out to get (1/6)*(-6) = -6/6 Note: you lose points because 6 is NOT prime --------------------------------------------------------------

jimthompson5910 (jim_thompson5910):

Let me know when you want me to continue

OpenStudy (anonymous):

You can continue

jimthompson5910 (jim_thompson5910):

ok we got these results -1/6, 2/6, 3/6, -4/6, 5/6, -6/6

jimthompson5910 (jim_thompson5910):

Add them up (-1/6) + (2/6) + (3/6) + (-4/6) + (5/6) + (-6/6) (-1+2+3+(-4) + 5 + (-6))/6 (-1+2+3-4+5-6)/6 -1/6

jimthompson5910 (jim_thompson5910):

So you will expect on average to lose 1/6 of a point every time you roll the die

OpenStudy (anonymous):

Oh! Okay thank you!

jimthompson5910 (jim_thompson5910):

you're welcome